### v1

@@ 5.2 Dataset
Table 2: Dataset used for evaluation. List of app categories: Productivity (PR), Communication (CO), Tools (TO), Social (SO), Health and fitness (HE), Personalization (PE), Travel and local (TR), Maps and Navigation (MA), Lifestyle (LI), Weather (WE).
@@ 5.3.1 Baseline fine-tuning
Table 3: Token classification results (baseline fine-tuning)
@@
Table 5: Human evaluation of new features (FP) with best-performing T-FREX baseline model (XLNetlarge). Total column is weighted based on review distribution across categories.
@@ 5.3.2 Extended pre-training
Figure 9 illustrates the evolution of all functional correctness metrics for the in-domain fine-tuning using model checkpoints $c$ from $1\rightarrow 10$. For a comparative analysis, we include T-FREX baseline model as model checkpoint $c=0$. Results showcase that all six model instances increase the maximum value for each metric at some point during the extended pre-training. Precision is on average the most increased metric, especially for base models BERTbase ($c=2$, +0.159), RoBERTabase ($c=5$, +0.166) and XLNetbase ($c=2$, +0.097). The only exception is precision for XLNetlarge, which suffers from decay from 0.761 to 0.686 (-0.075). The increase of recall is more modest and is especially highlighted in base models like XLNetbase ($c=4$, +0.078) but also in large models such as RoBERTalarge ($c=9$, +0.040) or XLNetlarge ($c=1$, +0.047).
(a) Results for BERTbase
(b) Results for BERTlarge
(c) Results for RoBERTabase
(d) Results for RoBERTalarge
(e) Results for XLNetbase
(f) Results for XLNetlarge
Figure 9: Token classification evaluation results with extended pre-training over different checkpoints (train epochs).
@@ 5.3.3 Instance selection
Figure 10 illustrates the evolution of all functional correctness metrics for the in-domain fine-tuning using T-FREX baseline approach for fine-tuning and the instance selection algorithm to generate different training data set partitions $d\in\{12.5\%,25\%,50\%,75\%\}$. For a comparative, evolutionary analysis, we include T-FREX baseline setting using the complete training set as $d=100\%$.
(a) Results for BERTbase
@@
(c) Results for RoBERTabase
(d) Results for RoBERTalarge
(e) Results for XLNetbase
(f) Results for XLNetlarge
Figure 10: Token classification evaluation results with instance selection over different training data partitions (of review instances).
@@ 5.3.4 Combining extended pre-training and instance selection
Table 6: Summary of best-performing approach according to each setting: baseline (BL), with extended pre-training (EP), with instance selection (IS), and combining extended pre-training with instance selection (EP/IS). Green cells represent the best metrics for each model. Pinpointed cells with [Uncaptioned image] icon represent the best metrics in all evaluations.
[…]
@@
EP c=10.738 [Uncaptioned image] c=10.667 [Uncaptioned image] c=10.700 [Uncaptioned image] c=10.677
@@
To better illustrate the contributions of our research, Table 7 reports metric variations for each T-FREX extension with respect to the T-FREX baseline design. On average, we observe that all base models, as well as mostly all large models, improve T-FREX baseline when extended pre-training and/or instance selection is applied. For large models, there are some minor exceptions like precision for XLNetlarge and f-measures for BERTlarge. Precision is the most increased metric, experiencing its greatest improvement when extended pre-training and instance selection are used in combination for BERTbase (+0.172). Recall is also increased, but its maximum improvement is more conservative, as observed with RoBERTalarge with combined extensions (+0.112) or in XLNetlarge with instance selection (+0.094).
Table 7: Delta variations $\delta$ for evaluation metrics with respect to T-FREX baseline (for best setting in each case).
[…]
BERTbase EP +0.159 +0.057 +0.089 +0.037
IS +0.142 +0.074 +0.100 +0.051
EP/IS +0.172 +0.061 +0.094 +0.041
BERTlarge EP +0.043 +0.050 +0.054 +0.012
IS +0.008 -0.014 -0.011 -0.050
EP/IS -0.027 0.005 -0.007 -0.036
RoBERTabase EP +0.166 +0.054 +0.083 +0.036
IS +0.071 +0.029 +0.041 +0.005
EP/IS +0.073 +0.023 +0.039 -0.001
RoBERTalarge EP +0.095 +0.046 +0.065 +0.009
IS +0.063 +0.103 +0.095 +0.058
EP/IS +0.081 +0.112 +0.093 +0.068
XLNetbase EP +0.097 +0.078 +0.079 +0.036
IS +0.065 +0.096 +0.082 +0.049
EP/IS +0.104 +0.076 +0.073 +0.035
XLNetlarge EP -0.075 +0.047 +0.002 -0.017
IS -0.023 +0.094 +0.054 +0.031
EP/IS -0.013 +0.068 +0.023 +0.009
@@
Table 8: Human evaluation of new features (FP) with best-performing T-FREX extended model (c=1XLNetlarge). Total column is weighted based on review distribution across categories.
@@ 6.1 Baseline fine-tuning
RQ1: The T-FREX baseline design effectively extracts feature mentions from mobile app reviews, enhancing the performance of syntactic-based baseline methods in both out-of-domain and in-domain learning settings. This improvement is observed in terms of both precision and recall.
For the most challenging scenario (i.e., out-of-domain), XLNetbase demonstrated the strongest performance, achieving a weighted f-measure ($f_{\beta}$) of 0.499. This model is particularly effective in reducing missed features (false negatives). However, XLNetlarge excelled in precision, achieving the highest score of 0.651, indicating its ability to minimize false features (false positives). Conversely, RoBERTalarge showed the weakest performance, highlighting its limited generalization capability for out-of-domain scenarios.
@@
Overall, the results highlight the effectiveness of T-FREX models, particularly the XLNet variants, in various evaluation contexts. n encoder-only LLM approach showcases better adaptation to the concept of a feature as used in an industrial setting like AlternativeTo, suppressing the limitations of deterministic approaches conditioned by the use of specific syntactic patterns that might not always generalize. Hence, we argue that T-FREX provides an effective software-based mechanism to reduce missing features in automatic review-based opinion mining pipelines used by practitioners and researchers.
@@ 6.2 Extended pre-training
RQ2: Extended pre-training consistently improves the precision and recall of feature extraction for all encoder-only models analysed in this research. Different models entail different outcomes, requiring different training epochs according to the size and pre-training task of the model to reach maximum improvement without forcing the decay of quality metrics.
Base models such as BERTbase, RoBERTabase, and XLNetbase demonstrate substantial improvements in performance metrics with extended pre-training, with precision showing the most significant increases. This suggests that base models have greater potential for refinement, and extended pre-training effectively reduces false positives, enhancing the likelihood that identified features are accurate. Similarly, while large models tend to also improve with extended pre-training, they also tend to experience a decay after the first few epochs, indicating a limit to the benefits of extended pre-training for these more complex models. This is also supported by the evolution of the evaluation loss, which is significantly limited after a few epochs of extended pre-training.
@@
On average, the best performance improvements are achieved after just one or two epochs of extended pre-training. Beyond this, additional epochs do not provide substantial benefits and, in some cases, lead to performance decay. For similar tasks and settings, this suggests that investing in extensive pre-training beyond a few epochs may not be adequate in terms of cost-effectiveness balance, given the diminishing returns and significant computational resources required. Researchers and practitioners conducting similar experiments (e.g., NER fine-tuning with encoder-only LLMs) should consider these factors when designing their training protocols to ensure efficient and effective use of computational resources.
@@ 6.3 Instance selection
RQ3: On average, instance selection not only improves the performance efficiency of the fine-tuning process but also leads to a significant improvement in the functional correctness of the token classification task, both in precision and recall. This improvement is especially observed for data partitions between 50%-75%, and it is consistently observed in base models, while large models present some exceptions.
As illustrated with the analysis of base models, optimal performance is achieved with different data partitions from the original set of reviews (e.g., 75% for BERTbase, 25% for RoBERTabase, and 50% for XLNetbase). This underscores that base models may struggle with redundancies and noise in the full dataset, and a more targeted data selection can enhance their performance. Large models, while also benefiting from instance selection, show some variability, particularly with XLNetlarge, which exhibits erratic behaviour, suggesting that its architecture might be more sensitive to the training set size and content. Interestingly, the performance across all models tends to peak at certain data partition sizes and then decline, indicating a local optimum. Overall, empirical results showcase the validity of the instance selection algorithm (Algorithm 2) proposed for token-level fine-tuning in the context of mobile app review feature extraction.
In addition to effectiveness, using a smaller, optimal portion of the dataset significantly reduces execution times. Training with 50% of the dataset, for instance, nearly halves the execution time while still improving performance on average. For even smaller partitions, such as 12.5%, the speedup can be as much as four to five times faster, which is particularly advantageous for large-scale applications where computational resources and time are critical. This efficiency gain, combined with improved model performance, makes instance selection a highly valuable strategy in real-world scenarios, where processing large batches of reviews for multiple tasks (e.g., sentiment analysis, content classification, feature extraction) becomes computationally expensive.
@@ 6.4 Combining extended pre-training and instance selection
When the precision of feature proposals is the priority, RoBERTabase stands out as the best model, especially for tasks where false positives are more tolerable than missing actual features. Conversely, for applications where recall is critical, XLNetlarge is superior, as missing a feature mention is more costly than incorrectly identifying one. This underscores the need for decision-making based on specific application scenarios, which is why detailed reporting of these metrics is essential. However, in the context of RE hairy tasks (Berry, 2021), we argue that high recall is more important than high precision, which is reflected in the weighting of $f_{\beta}$.
On average, base models benefit more from instance selection, particularly in terms of recall, demonstrating their sensitivity to data quality and relevance. On the other hand, large models show significant improvements with extended pre-training, whether used alone or in conjunction with instance selection. Notably, base models always experience improvements with some form of extension. This motivates the need for a cost-effectiveness analysis considering both absolute performance (Table 6) and relative improvements (Table 7) to determine the practicality of maintaining base versus large models.
Regarding new, unseen features, human evaluation indicates a substantial increase in the number of identified features with extended T-FREX design. Due to the large number of feature candidates in the dataset (i.e., tokens), even a slight improvement in recall ($+0.047$ with c=1XLNetlarge) leads to a significant increase in potential new features in absolute numbers (1,956 vs. 11,120). Consequently, despite a slight drop in the precision of new features ($-0.017$), the much higher number of new feature mentions identified suggests a lower likelihood of missing important features. This trade-off supports a more exhaustive feature extraction process, as previously argued.
@@ 6.5 Threats to validity
For the elicitation of threats to validity, we rely on the taxonomy proposed by Wohlin (2014). Concerning construct validity, we rely on functional correctness metrics based on token classification performance at the token level. However, the feature extraction process is not evaluated in a real-world setting, where feature predictions are used for a particular purpose. Due to the extensiveness of the empirical evaluation showcased in this research, we limited the scope of analysis to the feature extraction task, leading to future work on additional quality characteristics from ISO 25010, such as functional appropriateness or interaction capability of T-FREX as a component in a real mobile app review mining pipeline. In addition, the selection of $f_{\beta}$ is conditioned by the human evaluation results. Different experiments might lead to different $\beta$ values. To reduce the threats imposed by this, we reported extended results including precision, recall and harmonized $f_{1}$ in addition to $f_{\beta}$ to facilitate replication in other contexts. Furthermore, we conducted a human evaluation with external annotators using a large set of reviews and multiple annotators (5 per task), increasing the reliability of results by using average values.
@@ 8 Conclusions and future work
As future work, we plan on evaluating the generalization of T-FREX models to other mobile app domains. Specifically, we want to explore its ability to generalize and extract features in emerging, disruptive domains, such as AI-based applications. Furthermore, we plan to explore generalization beyond the scope of mobile applications, such as desktop applications or APIs, and other user-generated documents, such as issues and bug reports. In conclusion, we envisage that both the methodological contributions as well as the set of T-FREX LLM instances (fine-tuned and with extended pre-training), which are publicly available, might assist future researchers and practitioners by integrating these models into their own review-based NLP pipelines for opinion mining and decision-making tasks where features are considered a core descriptor.
Acknowledgements.

### v2

@@ 5.2 Dataset
Table 2: Evaluation dataset. List of app categories: Productivity (PR), Communication (CO), Tools (TO), Social (SO), Health and fitness (HE), Personalization (PE), Travel and local (TR), Maps and Navigation (MA), Lifestyle (LI), Weather (WE).
@@ 5.3.1 Baseline fine-tuning
Table 3: Token classification results (T-FREX baseline)
@@
Table 5: Human evaluation of new features (FP) with best-performing T-FREX baseline model (XLNetlarge).
@@ 5.3.2 Extended pre-training
Figure 9 illustrates the evolution of precision and recall metrics for the in-domain fine-tuning using model checkpoints $c$ from $1\rightarrow 10$. For a comparative analysis, we include T-FREX baseline model as model checkpoint $c=0$. Results showcase that all six model instances increase the maximum value for each metric at some point during the extended pre-training. Precision is on average the most increased metric, especially for base models BERTbase ($c=2$, +0.159), RoBERTabase ($c=5$, +0.166) and XLNetbase ($c=2$, +0.097). The only exception is precision for XLNetlarge, which suffers from decay from 0.761 to 0.686 (-0.075). The increase of recall is more modest and is especially highlighted in base models like XLNetbase ($c=4$, +0.078) but also in large models such as RoBERTalarge ($c=9$, +0.040) or XLNetlarge ($c=1$, +0.047).
(a) Base models
(b) Large models
Figure 9: Precision ($p$) and recall ($r$) evolution with extended pre-training.
@@ 5.3.3 Instance selection
Figure 10 illustrates the evolution of precision and recall metrics for the in-domain fine-tuning using T-FREX baseline approach for fine-tuning and the instance selection algorithm to generate different training data set partitions $d\in\{12.5\%,25\%,50\%,75\%\}$. For a comparative, evolutionary analysis, we include T-FREX baseline setting using the complete training set as $d=100\%$.
(a) Base models
@@
Figure 10: Precision ($p$) and recall ($r$) evolution with instance selection.
@@ 5.3.4 Combining extended pre-training and instance selection
Table 6: Metrics for the best-performing approach with background colors representing $\delta$ variations relative to the T-FREX baseline (BL). Bolded cells indicate the best metrics for each model. Pinpointed cells [Uncaptioned image] denote the overall best metrics across all evaluations.
model setting $p$ $r$ $f_{1}$ $f\textsubscript{$\beta$}$
@@
EP c=10.738 c=1[Uncaptioned image] 0.667 c=1[Uncaptioned image] 0.700 c=1[Uncaptioned image] 0.677
@@
Focusing on delta variations, we observe that on average all base models, as well as mostly all large models, improve T-FREX baseline when extended pre-training and/or instance selection is applied. For large models, there are some minor exceptions like precision for XLNetlarge and f-measures for BERTlarge. Precision is the most increased metric, experiencing its greatest improvement when extended pre-training and instance selection are used in combination for BERTbase (+0.172). Recall is also increased, but its maximum improvement is more conservative, as observed with RoBERTalarge with combined extensions (+0.112) or in XLNetlarge with instance selection (+0.094).
@@
Table 7: Human evaluation of new features (FP) with best-performing T-FREX extended model (c=1XLNetlarge).
@@ 6.1 Baseline fine-tuning
RQ1: The T-FREX baseline design improves upon syntactic-based baseline methods for extracting feature mentions from mobile app reviews (H1). These improvements are evident in both out-of-domain and in-domain learning settings, with notable gains in precision and recall.
For the most challenging scenario (i.e., out-of-domain), XLNetbase demonstrated the strongest performance, achieving a weighted f-measure ($f_{\beta}$) of 0.499, aligning with H1 even in less familiar contexts. This model is particularly effective in reducing missed features (false negatives). However, XLNetlarge excelled in precision, achieving the highest score of 0.651, indicating its ability to minimize false features (false positives). Conversely, RoBERTalarge showed the weakest performance, highlighting its limited generalization capability for out-of-domain scenarios.
@@
Overall, the results highlight the improved effectiveness of T-FREX models, particularly the XLNet variants, in various evaluation contexts, reinforcing H1 as they outperform traditional syntactic-based methods. An encoder-only LLM approach showcases better adaptation to the concept of a feature as used in an industrial setting like AlternativeTo, suppressing the limitations of deterministic approaches conditioned by the use of specific syntactic patterns that might not always generalize. Hence, we argue that T-FREX provides an effective software-based mechanism to reduce missing features in automatic review-based opinion mining pipelines used by practitioners and researchers.
@@ 6.2 Extended pre-training
RQ2: Extended pre-training consistently improves the precision and recall of feature extraction for all encoder-only models analysed in this research (H2). Different models entail different outcomes, requiring different training epochs according to the size and pre-training task of the model to reach maximum improvement without forcing the decay of quality metrics.
Base models demonstrate substantial improvements in performance metrics with extended pre-training, with precision showing the most significant increases. This suggests that base models have greater potential for refinement, and extended pre-training effectively reduces false positives, enhancing the likelihood that identified features are accurate. Similarly, while large models tend to also improve with extended pre-training, they also tend to experience a decay after the first few epochs, indicating a limit to the benefits of extended pre-training for these more complex models. This is also supported by the evolution of the evaluation loss, which is significantly limited after a few epochs of extended pre-training.
@@
On average, the best performance improvements are achieved after one or two epochs of extended pre-training, further validating H2 by showing that even limited extended pre-training enhances LLMs’ reflection of mobile app review contexts. Beyond this point, we observe similar or even diminishing results. In similar contexts, this suggests that investing in extensive pre-training beyond a few epochs may not be adequate in terms of cost-effectiveness balance, given the diminishing returns and significant computational resources required. Researchers and practitioners conducting similar experiments (e.g., NER fine-tuning with encoder-only LLMs) should consider these factors when designing their training protocols to ensure efficient and effective use of computational resources.
@@ 6.3 Instance selection
RQ3: On average, instance selection not only improves the performance efficiency of the fine-tuning process but also leads to an improvement in the functional correctness of the token classification task, both in precision and recall (H3). This improvement is especially observed for data partitions between 50%-75%, and it is consistently observed in base models, while large models present some exceptions.
As illustrated with the analysis of base models, optimal performance is achieved with different data partitions from the original set of reviews (e.g., 75% for BERTbase, 25% for RoBERTabase, and 50% for XLNetbase). This underscores that base models may struggle with redundancies and noise in the full dataset, and a more targeted data selection can enhance their performance. Large models, while also benefiting from instance selection, show some variability, particularly with XLNetlarge, which exhibits erratic behaviour, suggesting that its architecture might be more sensitive to the training set size and content. Interestingly, the performance across all models tends to peak at certain data partition sizes and then decline, indicating a local optimum. Results showcase the validity of the instance selection method proposed for token-level fine-tuning in the context of mobile app review feature extraction.
In addition to effectiveness, using a smaller, optimal portion of the dataset significantly reduces execution times. Training with 50% of the dataset nearly halves the execution time while still improving performance on average. For even smaller partitions, such as 12.5%, the speedup can be as much as four to five times faster, which is particularly advantageous for large-scale applications where computational resources and time are critical. This efficiency gain, combined with improved model performance, makes instance selection a highly valuable strategy in real-world scenarios, where processing large batches of reviews for multiple tasks (e.g., sentiment analysis, content classification, feature extraction) becomes computationally expensive.
@@ 6.4 Combining extended pre-training and instance selection
When focusing on precision, RoBERTabase stands out as the best model, especially for tasks where false positives are more tolerable than missing actual features. Conversely, for applications where recall is critical, XLNetlarge is superior, as missing a feature mention is more costly than incorrectly identifying one. This underscores the need for decision-making based on specific application scenarios, which is why detailed reporting of these metrics is essential. However, in the context of RE hairy tasks (Berry, 2021), we argue that high recall is more important than high precision, which is reflected in the weighting of $f_{\beta}$.
On average, base models benefit more from instance selection, particularly in terms of recall, demonstrating their sensitivity to data quality and relevance. On the other hand, large models show significant improvements with extended pre-training, whether used alone or in conjunction with instance selection. Notably, base models always experience improvements with some form of extension. This motivates the need for a cost-effectiveness analysis considering both absolute performance and relative improvements (Table 6) to determine the practicality of maintaining base versus large models.
Regarding new, unseen features, human evaluation indicates a substantial increase in the number of identified features with extended T-FREX design. Due to the large number of feature candidates in the dataset (i.e., tokens), a slight improvement in recall ($+0.047$, c=1XLNetlarge) leads to a significant increase in potential new features in absolute numbers (1,956 vs. 11,120). Consequently, despite a slight drop in the precision of new features ($-0.017$), the much higher number of new feature mentions identified suggests a lower likelihood of missing important features. This trade-off supports a more exhaustive feature extraction process, as previously argued.
@@ 6.5 Threats to validity
For the elicitation of threats to validity, we rely on the taxonomy proposed by Wohlin (2014). Concerning construct validity, we rely on functional correctness metrics based on token classification performance at the token level. However, the feature extraction process is not evaluated in a real-world setting, where feature predictions are used for a particular purpose. Due to the extensiveness of the empirical evaluation showcased in this research, we limited the scope of analysis to the feature extraction task, leading to future work on additional quality characteristics from ISO 25010, such as functional appropriateness or interaction capability of T-FREX as a component in a real mobile app review mining pipeline. In addition, the selection of $f_{\beta}$ is conditioned by the human evaluation results. Different experiments might lead to different $\beta$ values. To reduce the threats imposed by this, we reported extended results including precision, recall and harmonized $f_{1}$ in addition to $f_{\beta}$ to facilitate replication in other contexts. Furthermore, we conducted a human evaluation with external annotators using a large set of reviews and multiple annotators (5 per task), increasing the reliability of results by using average values. Finally, our human evaluation uses non-developer participants, which may not fully reflect real-world expertise. To address this, we include precision and recall metrics alongside $f_{\beta}$ in our analysis for a balanced evaluation.
@@ 8 Conclusions and future work
As future work, we plan on evaluating the generalization of T-FREX models to other mobile app domains. Specifically, we want to explore its ability to generalize and extract features in emerging, disruptive domains, such as AI-based applications. Furthermore, we plan to explore generalization beyond the scope of mobile applications, such as desktop applications or APIs, and other user-generated documents, such as issues and bug reports. Finally, the functional appropriateness and suitability of T-FREX in a real-world setting remain to be explored, particularly its adoption within review mining pipelines for user feedback analysis. In conclusion, we envisage that both the methodological contributions as well as the set of T-FREX LLM instances (fine-tuned and with extended pre-training), which are publicly available, might assist future researchers and practitioners by integrating these models into their own review-based NLP pipelines for opinion mining and decision-making tasks where features are considered a core descriptor.
@@
Funding
Ethical Approval
This study adhered to established ethical principles for research involving human participants. As data collection was fully anonymous and conducted through online platforms, formal ethical approval was not required.
Informed Consent
All participants provided informed consent before participating in the study. They were informed about the research purpose, data usage, and their right to withdraw at any time. No personally identifiable information was collected.
Author Contributions
Quim Motger: Conceptualization, Methodology, Software, Validation, Formal analysis, Investigation, Data Curation, Writing - Original Draft, Visualization. Alessio Miaschi: Conceptualization, Methodology, Validation, Investigation, Resources, Writing - Original Draft. Felice Dell’Orletta: Conceptualization, Methodology, Writing - Review & Editing, Supervision, Funding Acquisition. Xavier Franch: Conceptualization, Writing - Review & Editing, Supervision, Funding Acquisition. Jordi Marco: Conceptualization, Writing - Review & Editing, Supervision.